OpenStudy (anonymous):

Write the equation in standard or vertex form of y=x^2+5x+4 with the vertex of (-2.5 , -2.25)

OpenStudy (ranga):

Complete the square of (x^2+5x+4) and you will have it in the vertex form.

OpenStudy (anonymous):

thank you

OpenStudy (ranga):

you are welcome. I assume you know how to complete the square.

OpenStudy (anonymous):

I have an idea of what it is but I am not sure sure if it is what I'm thinking

OpenStudy (ranga):

x^2+5x+4 Take the coefficient of the x term and divide it by 2: Here it is 5/2. This is the number that will go inside the square: (x + 5/2)^2. If you expand that you will get: x^2 + 5x + (5/2)^2 Since we are introducing the extra (5/2)^2 we need to subtract (5/2)^2 So x^2+5x+4 = (x + 5/2)^2 - (5/2)^2 + 4 = (x + 5/2)^2 - 25/4 + 4 = (x + 5/2)^2 - 25/4 + 16/4 = (x + 5/2)^2 + (-25 + 16) / 4 = (x + 5/2)^2 - 9/4 = (x + 2.5)^2 - 2.25 = (x + 2.5)^2 - 2.25 = (x + 2.5)^2 - 2.25 x^2+5x+4 = (x - (-2.5))^2 + (-2.25) y = (x - (-2.5))^2 + (-2.25) is the vertex form where the vertex is (-2.5, -2.25)

OpenStudy (ranga):

The general vertex form of a parabola is: y = a(x - h)^2 + k where (h,k) is the vertex.

OpenStudy (anonymous):

Thank you so much. I really do appreciate it :)

OpenStudy (ranga):

you are very welcome.